# GST: Precise 3D Human Body from a Single Image with Gaussian Splatting Transformers

### v1

@@ Abstract
Reconstructing realistic 3D human models from monocular images has significant applications in creative industries, human-computer interfaces, and healthcare. We base our work on 3D Gaussian Splatting (3DGS), a scene representation composed of a mixture of Gaussians. Predicting such mixtures for a human from a single input image is challenging, as it is a non-uniform density (with a many-to-one relationship with input pixels) with strict physical constraints. At the same time, it needs to be flexible to accommodate a variety of clothes and poses. Our key observation is that the vertices of standardized human meshes (such as SMPL) can provide an adequate density and approximate initial position for Gaussians. We can then train a transformer model to jointly predict comparatively small adjustments to these positions, as well as the other Gaussians’ attributes and the SMPL parameters. We show empirically that this combination (using only multi-view supervision) can achieve fast inference of 3D human models from a single image without test-time optimization, expensive diffusion models, or 3D points supervision. We also show that it can improve 3D pose estimation by better fitting human models that account for clothes and other variations. The code is available on the project website https://abdullahamdi.com/gst/.
Figure 1: Gaussian Splatting Transformers. We propose a method that, given a single input image, predicts an accurate 3D human pose and shape, along with a color model that enables novel view rendering, including clothing. GST relies solely on multi-view supervision (no direct 3D supervision).
@@ 1 Introduction
Ultraman [7] 20 minutes Optimization Mesh
ANIM [44] few seconds Inference SDF Direct 3D
@@
Reconstructing realistic 3D human models from monocular images is crucial for virtual reality and creative industries, as well as possibly improving human pose estimation for human-computer interfaces and health applications. It is also an integral component of “3D spatial computing” for mainstream consumer products incorporating 3D vision in VR and augmented reality (AR). To be practical in homes, offices, and workplaces, these products require precise 3D rendering, speed, compactness, flexibility, and realism. However, reconstructing 3D models from a single image remains challenging. Previous approaches that have shown success in this problem often utilize 2D priors in a multi-view setup [14, 45, 32, 8, 38, 36, 46, 20, 47, 40, 57]. Specifically, in human modeling, issues arise due to intricate 3D details such as facial features, clothing, and joints, which present challenges for deep learning methods. Early methods addressed these challenges using a learned Signed Distance Function (SDF) on a human template to predict detailed 3D meshes [53, 70]. Later works incorporated Neural Radiance Fields (NeRFs) to capture texture details [24, 61], or leveraged pre-trained diffusion models to generate dense views from a single frontal image, reducing prediction ambiguity [61, 22, 71, 65, 16, 7]. However, they typically suffer from low speed, hindering real-time deployment.
@@
In this work, we present GST (Gaussian Splatting Transformer), illustrated in Fig. 1, a direct method that learns to predict 3D Gaussian Splatting [28] for 3D representation, allowing for fast rendering and flexible editing abilities compared to others. Our method does not rely on diffusion priors and is, therefore, capable of near real-time predictions. This is essential for downstream applications and ensures that the inference can be easily incorporated with prior-based approaches. GST leverages multi-view supervision instead of precise (and expensive) 3D point clouds. Despite this, it predicts accurate 3D joint and body poses while maintaining the perceptual quality of renderings from novel views. Table 1 summarizes the characteristics of prior work.
@@
GST is inspired by recent works on single view 3D reconstruction [56]. However, the complexity of human pose in 3D space poses significant challenges to the direct applications of methods that associate one 3D point (or Gaussian) to each pixel. Therefore, we augment our model also to predict the pose parameters of the SMPL [37] model. The SMPL model is used as the scaffolding on which the Gaussians are positioned and rendered. Each Gaussian is loosely tied to a vertex on the SMPL model by an offset. This has two advantages. First, it provides a good initialization of the density and pose of the Gaussians, including back faces, which are notoriously difficult for single-view methods. Second, we find that the joint optimization of pose and appearance also improves the SMPL pose prediction.
@@
To the best of our knowledge, GST is the first work that efficiently combines accurate 3D human prediction with improved visual quality, utilizing only multi-view supervision and without relying on diffusion priors. In summary, our contributions are the following:
@@
1) We propose GST, a 3D human body model prediction method that does not rely on diffusion priors and performs novel view synthesis from a single image input. This makes it particularly amenable to real-time modeling tasks, where multiple views are uneconomical or impractical.
2) We evaluate our method and compare it to other state-of-the-art models. Although prior methods only solve for 3D pose estimation or 3D reconstruction, our method still performs equally or better on perceptual and 3D pose estimation metrics without 3D supervision.
@@ 2 Related work
3D Reconstruction using Image Priors. In the area of prior-based 3D reconstruction, contemporary zero-shot text-to-image generators [48, 51, 49, 52, 2, 14] have shown significant improvement by leveraging enhanced synthesis priors [45, 60, 39, 6, 9]. DreamFusion [45] stands out as a pioneering work that distilled a pre-existing diffusion model [52] into a NeRF framework [41, 3] using text prompts. This innovation spurred further research in both text-to-3D synthesis [32, 8] and image-to-3D reconstruction [54, 38, 36, 46]. The latter approach leverages supplementary reconstruction losses focused on frontal camera perspectives [36] and subject-specific diffusion guidance [47, 46]. In addition, task-specific priors have been explored [23, 26, 50], as well as additional control mechanisms [40]. More recently, Gaussian-Splatting approaches [28, 21] have improved the efficiency of 3D generation optimization through rapid Gaussian Splatting rasterization  [57, 58, 68]. In contrast, our method GST does not rely on diffusion priors, providing a simpler and more cost-effective framework specialized for human 3D reconstruction and jointly predicts the precise internal 3D human body joints.
Figure 2: Overview of the pipeline of Gaussian Splatting Transformer (GST). Given a single input image, GST uses a Vision Transformer (ViT) to predict both a 3D human pose (in the form of SMPL parameters) and a refined full-color 3D model (in the form of 3D Gaussian Splats). Additional input tokens facilitate the output of individual Gaussians’ color $\mathbf{c}$, opacity $\alpha$, scale, rotation, and position offset $\bm{\delta}$. Each Gaussian’s position $\bm{\mu}$ is relative to one vertex of the SMPL model $\mathbf{v}$ by the offset $\bm{\delta}$, and so this model can be considered a refinement or residual over the interpretable SMPL mesh, facilitating multi-view rendering with higher visual fidelity.
@@
3D Human Pose Estimation. Many approaches in the literature focus on predicting 3D human pose and shape from a single image [27, 30, 33, 10, 34, 18]. Relevant for our work are the approaches that directly regress the body shape and pose from a single image. The first work to introduce this approach was HMR [27], which uses a CNN to regress SMPL [37] parameters. Dedicated designs have been proposed for the HMR architecture; HoloPose [19] suggests a pooling strategy based on the 2D locations of body joints, while HKMR [17] relies on SMPL hierarchical structure to make predictions. PARE [29] introduces a body-part-guided attention mechanism to handle occlusions better, and PyMAF [66, 67] incorporates a mesh alignment module for SMPL parameter regression. More recently HMR2 [18] utilizes a transformer to predict the SMPL parameters and train on a large pool of 3D data and unprotected 2D joint labels, while TokenHMR [12] improved HMR2 by leveraging tokenized encoding and reduced the 2D basis in training HMR2. In contrast to all of these methods, our GST does not rely on 3D supervision and utilizes novel view synthesis training of a transformer to predict the Gaussian splats that are grounded on predicted SMPL parameters.
@@
Monocular 3D Human Reconstruction. Recent advancements in 3D human reconstruction from single images have resulted in diverse methods, each employing different data sources, 3D representations, and supervision strategies [61, 24, 22, 71, 70, 65, 13, 16, 7, 44, 64, 25]. PIFU [53] is one of the earlier works that successfully uses the learned Sign Distance Function (SDF) representation with direct 3D supervision to reconstruct a detailed 3D mesh of humans from a single image. ANIM [44] incorporates sparse voxel depth features with the input image features and uses direct 3D supervision to train on the RGBD input (depth is needed). SHERF [24] developed on PIFU’s 3D representation and adopted a NeRF representation for decoding the 3D human, training with multi-view supervision. While GST follows SHERF in the multi-view supervision, we utilize the more explicit Gaussian Splatting representation [28] for 3D, allowing for more flexible control and better 3D alignment. Similar to our method, A-NeRF [55] jointly optimizes human pose and 3D reconstruction, however it takes videos as input and does not generalise to unseen subjects.
@@
With the recent wave of success of generative image and text models  [48, 51, 49, 52, 42], several methods try to leverage these foundation models to improve the performance of 3D human reconstruction [71, 22, 61, 13, 7, 16]. For example, SIFU [71] integrates GPT-predicted captions with diffusion models for back-view generation and texture refinement and builds on GTA [70], its predecessor, which learns a triplane SDF decoder. Similarly, SiTH [22] employees Diffusion-prior to generate back views and decode the SDF and texture colours, while HumanLRM [61] generate multi-views with pre-trained diffusion and then train a tri-plane NeRF Large Reconstruction Model (LRM) for decoding. SHERT [65] utilizes semantic mesh and whole texture inpainting with the help of diffusion priors to create detailed 3D Mesh of humans. Most of these methods use pre-trained diffusions to improve the texturing with optimization, which slows down the process and prevents scaling for long videos, unlike our GST, which does not use any diffusion priors and runs at almost real-time inference, allowing for flexibility and potential integration with priors.
@@ 3.1 Architecture
Our model predicts 3D Gaussian splatting parameters from a single input image using a transformer architecture, including tokenization, processing through transformer blocks, and decoding into Gaussian parameters. We detail the model architecture (3.1) and the loss functions (3.2).
@@
This representation could be learned with image-pose pairs $(\mathbf{X},\bm{\theta},\bm{\beta})$. However, here we focus on multi-view supervision, as 3D supervision is costly and scarce.
@@
This combination ensures the 3D model captures both shape and appearance, allowing more realistic reconstructions.
@@ 3.2 Loss Functions
where $\lambda_{\text{tight}}$ is the weighting hyperparameter for the tightness regularization. As we show later in Sec. 5.3, this regularization plays an important role in the precision of the 3D human body predicted by GST. By minimizing this combined loss, our GST model learns to generate accurate and visually pleasing 3D reconstructions from a single image.
       Input Image        GT View        Out View        GT        Ours        HMR2 [18]        SMPL Overlays
Figure 3: 3D SMPL Shape. 3D human body results of our GST on two subjects of HuMMan [4] dataset compared to Ground Truth renderings, Ground Truth SMPL parameters [37], and SMPL predictions of HMR2 [18]. The overlay of the 3 SMPL bodies (ours, HMR2, and GT) shows that our predicted SMPL is more precise, while our predicted Gaussian splats maintain visual quality from novel views.
@@ 4 Experiments
In this section we describe our evaluation setup and the baselines for our comparisons.
@@ 4.1 Datasets and Metrics
Datasets. Similar to previous works [24], we utilize four comprehensive human datasets for evaluation: THuman [72], RenderPeople [1], ZJU MoCap [43], and HuMMan [4]. For ZJU MoCap, the dataset is divided following the SHERF setup [24]. Similarly, for HuMMan, we adhere to the official split (HuMMan-Recon), using 317 sequences for training and 22 for testing, with 17 frames sampled per sequence. For THuman, we select 90 subjects for training and 10 for testing, and for the RenderPeople dataset, we randomly sample 450 subjects for training and 30 for testing. Those four datasets used for evaluation above are all small in terms of subject diversity. To showcase the capabilities of GST on a large dataset, we train our GST also on the TH21 dataset [63], which contains 2,500 3D scans, with diverse subject diversity. We randomly select 200 scans for evaluation.
Table 2: Human Novel View Synthesis and 3D Keypoints Evaluation Performance Comparison. We compare GST on the RenderPeople [1] and HuMMan [4] datasets. For each dataset, we report PSNR, SSIM, and LPIPS for novel view synthesis, as well as MPJPE (in mm) for 3D keypoints evaluation. The top section methods use the Ground Truth input SMPL parameters, which are shown for reference, while the bottom section methods only use the single image input (our setup). $\uparrow$ means the larger is better; $\downarrow$ means the smaller is better.
@@
Baselines. In addition to earlier works on Human NeRF with multi-view setting, NHP [31] and MPS-NeRF [15], we compare to recent single-image methods SHERF [24] for novel view synthesis and HMR2 [18] and TokenHMR [12] for 3D Human reconstruction precision. Different from SHERF, our method does not take as input ground truth SMPL parameters, therefore we adapt SHERF to use HMR2 [18] or TokenHMR [12] SMPL predictions for a fair comparison to our method. We also include a fast and salable Splatter Image [56], a state-of-the-art single image 3D reconstruction method for novel view synthesis tables.
@@ 4.2 Implementation Details of GST
Our model follows the implementation of HMR2 [18] for the predictions of SMPL parameters. We extend the HMR2 decoder implementation to process some additional learnable tokens for the predictions of the Gaussian parameters. The Gaussian parameters (color, rotation, scale, opacity, offset) are predicted for $K=26$ groups of 265 Gaussians. The token output is passed through a linear layer to obtain the final parameters. We use pre-trained weights from HMR2 for the ViT and the decoder and freeze the ViT weights during training. For our experiments, we use the loss weights $\mathcal{L}_{\text{perceptual}}=0.01$, $\mathcal{L}_{\alpha}=0.1$, $\mathcal{L}_{\text{tight}}=0.1$, and we train on square image crops of size 256. We train on a single A6000 GPU with a batch size of 32 for 3 days. At test time, GST can simultaneously perform 3D human pose estimation and 3D reconstruction in a single forward pass at 47fps.

### v2

@@ Abstract
Reconstructing posed 3D human models from monocular images has important applications in the sports industry, including performance tracking, injury prevention and virtual training. In this work, we combine 3D human pose and shape estimation with 3D Gaussian Splatting (3DGS), a representation of the scene composed of a mixture of Gaussians. This allows training or fine-tuning a human model predictor on multi-view images alone, without 3D ground truth. Predicting such mixtures for a human from a single input image is challenging due to self-occlusions and dependence on articulations, while also needing to retain enough flexibility to accommodate a variety of clothes and poses. Our key observation is that the vertices of standardized human meshes (such as SMPL) can provide an adequate spatial density and approximate initial position for the Gaussians. We can then train a transformer model to jointly predict comparatively small adjustments to these positions, as well as the other 3DGS attributes and the SMPL parameters. We show empirically that this combination (using only multi-view supervision) can achieve near real-time inference of 3D human models from a single image without expensive diffusion models or 3D points supervision, thus making it ideal for the sport industry at any level. More importantly, rendering is an effective auxiliary objective to refine 3D pose estimation by accounting for clothes and other geometric variations. The code is available at https://github.com/prosperolo/GST.
[Uncaptioned image]Input ImageOther ViewsHMR2 [12]GST (ours)Input ImageOther ViewsHMR2 [12]GST (ours)
Figure 1: Example of human pose improvements using our method GST. 3D human body results of our GST and SMPL predictions of HMR2 [12] on a sports sequence from the CMU panoptic dome dataset [20].
@@ 1 Introduction
Creating posed 3D human models from monocular images is crucial for the sports industry since reliably detecting the shape and pose of all humans in the scene is a fundamental prerequisite for any system monitoring players’ health or performance. A reliable system monitoring players’ poses throughout different games could be used to analyze their skills training, monitor return to play after injury, and influence game regulations to make the sport safer. These products require precise 3D rendering, speed, compactness, and flexibility to be practical and economically viable for teams at any level.
Popular approaches like HMR2 [12] regress the parameters of a human body model such as SMPL [29]. However, these approaches require ground truth 3D pose and shape annotations for every training frame. Therefore, they require significant time and economic investment for data collection before training in any new domain. This makes these methods unsuitable for training in any specific sports setup. Moreover, SMPL can only model the shape of the human body and cannot account for clothing or hair deformations.
@@
Modeling intricate 3D details and deformations of facial features, clothing, and hair is a long-standing challenge for deep learning methods. Early methods addressed these challenges using a learned Signed Distance Function (SDF) on a human template to predict detailed 3D meshes [39, 53]. Later works incorporated Neural Radiance Fields (NeRFs) to capture texture details [15, 43] or leveraged pre-trained diffusion models to generate dense views from a single frontal image, reducing prediction ambiguity [43, 14, 54, 48, 10, 4]. Despite the good visual quality of the results, these methods suffer from low speed, hindering real-time deployment, and often require expensive 3D scans as supervision.
@@
In this work, we present GST (Gaussian Splatting Transformer), illustrated in Fig. 2, a direct method that learns to predict 3D Gaussian Splatting [22] as 3D representation, allowing for fast rendering and flexible editing abilities compared to others. Our method does not rely on diffusion priors and is, therefore, capable of near real-time predictions. GST leverages multi-view supervision instead of precise (and expensive) 3D point clouds. Despite this, it predicts accurate 3D joint and body poses while maintaining the perceptual quality of renderings from novel views. Table 1 summarizes the characteristics of prior work.
@@
GST is inspired by recent works on single-view 3D reconstruction [41]. However, the complexity of human pose in 3D space poses significant challenges to the direct applications of methods that associate one 3D point (or Gaussian) to each pixel. Therefore, we also augment our model to predict the pose parameters of the SMPL [29] model. The SMPL model is used as the scaffolding on which the Gaussians are positioned and rendered. Each Gaussian is loosely tied to a vertex on the SMPL model by an offset. This has two advantages. First, it provides a good initialization of the density and pose of the Gaussians, including back faces, which are notoriously difficult for single-view methods. Second, we find that the joint optimization of pose and appearance improves the SMPL pose prediction.
@@
To the best of our knowledge, GST is the first work that efficiently combines fast and accurate 3D pose human prediction with improved geometry, utilizing only multi-view supervision and without relying on diffusion priors. In summary, our contributions are the following:
@@
Contributions: (i) We propose GST, a 3D human body model prediction method that does not rely on diffusion priors and performs novel view synthesis and human pose estimation from a single image input. This makes it particularly amenable to real-time modeling tasks, where multiple views are uneconomical or impractical. (ii) We evaluate our method and compare it to other state-of-the-art models. In contrast to prior methods that only solve for 3D pose estimation or 3D reconstruction, our method can predict both without 3D supervision, making it suitable for training and deployment on any specific sports setup.
Figure 2: Overview of the pipeline of Gaussian Splatting Transformer (GST). Given a single input image, GST uses a Vision Transformer (ViT) to predict both the 3D human pose (in the form of SMPL parameters) and a refined full-color 3D model (in the form of 3D Gaussian Splats). Additional input tokens are used to predict each Gaussian color $\mathbf{c}$, opacity $\alpha$, scale, rotation, and position offset $\bm{\delta}$. Every Gaussian position $\bm{\mu}$ is tied to one vertex of the SMPL model $\mathbf{v}$ by the offset $\bm{\delta}$.
@@ 2 Related work
3D Human Pose Estimation. Many approaches in the literature focus on predicting 3D human pose and shape from a single image [21, 24, 26, 5, 27, 12]. The approaches that directly regress the body shape and pose from a single image are most relevant to our work. The first work to introduce this approach was HMR [21], which uses a CNN to regress SMPL [29] parameters. Dedicated designs have been proposed for the HMR architecture; HoloPose [13] suggests a pooling strategy based on the 2D locations of body joints, while HKMR [11] relies on SMPL hierarchical structure to make predictions. PARE [23] introduces a body-part-guided attention mechanism to handle occlusions better, and PyMAF [49, 50] incorporates a mesh alignment module for SMPL parameter regression. More recently, HMR2 [12] utilizes a transformer to predict the SMPL parameters and train on a large pool of 3D data and unprotected 2D joint labels. In contrast, TokenHMR [7] improved HMR2 by leveraging tokenized encoding and reduced the 2D basis in training HMR2. In contrast to all these methods, our GST does not rely on 3D supervision and utilizes novel view synthesis training of a transformer to predict the Gaussian splats grounded on predicted SMPL parameters. Similar to our method, A-NeRF [40] jointly optimizes human pose and 3D reconstruction. However, it takes videos as input and does not generalise to unseen subjects.
@@
Single-image Human Reconstruction. Recent advancements in 3D human reconstruction from single images have resulted in diverse methods, each employing different data sources, 3D representations, and supervision strategies [43, 15, 14, 54, 53, 48, 8, 10, 4, 33, 47, 16]. PIFU [39] is one of the earlier works that successfully uses the learned Sign Distance Function (SDF) representation with direct 3D supervision to reconstruct a detailed 3D mesh of humans from a single image. ANIM [33] incorporates sparse voxel depth features with the input image features and uses direct 3D supervision to train on the RGBD input (depth is needed). SHERF [15] developed on PIFU’s 3D representation and adopted a NeRF representation for decoding the 3D human, training with multi-view supervision. While GST follows SHERF in the multi-view supervision, we utilize the more explicit Gaussian Splatting representation [22] for 3D, allowing for more flexible control and better 3D alignment.
@@
With the recent wave of success of generative image and text models  [35, 37, 36, 38, 30], several methods try to leverage these foundation models to improve the performance of 3D human reconstruction [54, 14, 43, 8, 4, 10]. For example, SIFU [54] integrates GPT-predicted captions with diffusion models for back-view generation and texture refinement and builds on GTA [53], its predecessor, which learns a triplane SDF decoder. Similarly, SiTH [14] employs a diffusion prior to generate back views and decode the SDF and texture colours, while HumanLRM [43] generates multi-views with a pre-trained diffusion model and then trains a tri-plane NeRF Large Reconstruction Model (LRM) for decoding. ELICIT [17] leverages CLIP [34] to generate text-conditioned unseen regions. SHERT [48] utilizes semantic mesh and whole texture inpainting with the help of diffusion priors to create detailed 3D meshes of humans. HumanSplat [31] and Human 3diffusion [44] use diffusion priors to generate novel views, and TeCH [18] uses text-to-image diffusion models to embed indescribable appearance information. Most of these methods rely on pre-trained diffusion models for texturing optimization, which slows down the process and limits scalability for high-speed, long-duration video motion. In contrast, our GST operates without diffusion priors, achieving near real-time inference while remaining flexible and capable of integrating with priors if needed.
@@ 3.1 Architecture
Our model predicts 3D Gaussian splatting parameters from a single input image using a transformer architecture, including tokenization, processing through transformer blocks, and decoding into Gaussian parameters. We detail the model architecture (Sec. 3.1) and the loss functions (Sec. 3.2).
@@
This representation could be learned with image-pose pairs $(\mathbf{X},\bm{\theta},\bm{\beta})$. However, here, we focus on multi-view supervision, as 3D supervision is costly and scarce, and not readily available for real-world setups.
@@
This combination ensures that the 3D model captures both shape and appearance, allowing for more realistic reconstructions.
@@ 3.2 Loss Functions
where $\lambda_{\text{tight}}$ is the weighting hyperparameter for the tightness regularization. As we show later in Sec. 5.3, this regularization plays an important role in the precision of the 3D human body predicted by GST. By minimizing this combined loss, our GST model learns to generate accurate human 3D models from a single image.
Optionally, a further regularization term for the SMPL shape parameters can also be included. This regularization term is an L2 loss on the $\beta$ parameters. This additional term does not affect the pose or visual metrics. It only influences the predicted body thickness.
@@ 4 Experiments
This section describes our evaluation setup and the baselines and prior work used in our comparisons.
@@ 4.1 Datasets and Metrics
Datasets. Similar to previous works [15], we utilize four comprehensive human datasets for evaluation: THuman [55], RenderPeople [1], ZJU MoCap [32], and HuMMan [2]. For ZJU MoCap, the dataset is divided following the SHERF setup [15]. Similarly, for HuMMan, we adhere to the official split (HuMMan-Recon), using 317 sequences for training and 22 for testing, with 17 frames sampled per sequence. For THuman, we select 90 subjects for training and 10 for testing, and for the RenderPeople dataset, we randomly sample 450 subjects for training and 30 for testing. Those four datasets used for evaluation above are all small in terms of subject diversity. To showcase the capabilities of GST on a large dataset, we train our GST also on the TH21 dataset [46], which contains 2,500 3D scans with high subject diversity. We randomly selected 200 scans for evaluation. Further validation is also conducted on the CMU Panoptic dataset with the single human partition that includes 9 sequences with 31 HD camera views; some qualitative examples of the pose estimation results for the sports sequence in this dataset are shown in Fig. 1 and in the Appendix. Another dataset that could have been suitable for evaluating our method is the SportsPose dataset [19]; however, no multi-view data has been released.
Table 2: Human Novel View Synthesis and 3D Keypoints Evaluation Performance Comparison. We compare GST on the RenderPeople [1] and HuMMan [2] datasets. We report PSNR, SSIM, and LPIPS for novel view synthesis and MPJPE (in mm) for 3D keypoint evaluation for each dataset. The top section methods use the Ground Truth input SMPL parameters, shown for reference, while the bottom section methods only use the single image input (our setup). $\uparrow$ means the larger is better; $\downarrow$ means the smaller is better. The best results are highlighted in bold.
@@
Baselines. In addition to earlier work on Human NeRF with a multi-view setting, NHP [25] and MPS-NeRF [9], we compare to recent single-image methods SHERF [15] for novel view synthesis and HMR2 [12] and TokenHMR [7] for 3D Human reconstruction precision. Unlike SHERF, our method does not take as input ground truth SMPL parameters. Therefore, we adapt SHERF to use HMR2 [12] or TokenHMR [7] SMPL predictions for a fair comparison to our method. We also include a comparison with Splatter Image [41], a state-of-the-art single-image 3D reconstruction method for novel-view synthesis tables.
Refer to captionRefer to captionInput ImageOther ViewsGTGST (ours)vs GTHMR2 [12]vs GT
Figure 3: 3D Shape Comparison with HMR2. 3D human body results of our GST on two subjects of HuMMan [2] dataset compared to Ground Truth SMPL parameters [29], and SMPL predictions of HMR2 [12].
@@ 4.2 Implementation Details of GST
Our model follows the implementation of HMR2 [12] for the prediction of the SMPL parameters. We extend the HMR2 decoder implementation to process some additional learnable tokens for the predictions of the Gaussian parameters. The Gaussian parameters (color, rotation, scale, opacity, and offset) are predicted for $K=26$ groups of 265 Gaussians. The token output is passed through a linear layer to obtain the final parameters. We use pre-trained HMR2 weights for the ViT and the decoder and freeze the ViT weights during training. For our experiments, we use the loss weights $\mathcal{L}_{\text{perceptual}}=0.01$, $\mathcal{L}_{\alpha}=0.1$, $\mathcal{L}_{\text{tight}}=0.1$, and we train on square image crops of size 256. We train on a single A6000 GPU with a batch size of 32 for 3 days. At test time, GST can simultaneously perform 3D human pose estimation and 3D reconstruction in a single forward pass at 47fps. We verify that the network can learn to predict high-quality renderings by overfitting on a single dataset example (cf. the Appendix for more details).